OpenStudy (anonymous):

Multiply out t(t^2-3)^2 & write the answer as a polynomial in t.

OpenStudy (anonymous):

howd you get this answer?

OpenStudy (saifoo.khan):

Give it a try..

OpenStudy (anonymous):

i would saifoo, if i knew how..grrr

OpenStudy (anonymous):

=t(t^4-6t^2+9) =t^5-6t^3+9t

OpenStudy (anonymous):

nalita- i see that, but im not sure how u got the answer. lol

OpenStudy (anonymous):

i simply get the product of the (t^2-3)^2,then i multiply it to t

OpenStudy (anonymous):

ok, i see now :) thank you.